OpenStudy (anonymous):

mn^2 (m +mny+ 3m^2) using distributive property. Please show steps to solve

OpenStudy (anonymous):

(mn)^2(1+y)+3(mn)^2=(3+y)(mn)^2=3(mn)^2+y(mn)^2